Determine the mass percent composition of a solution made by dissolving 5.31 moles CH3OH in 3.21 L of H,O. % CH3OH
Answer #1

Molar mass of CH3OH,

MM = 1*MM(C) + 4*MM(H) + 1*MM(O)

= 1*12.01 + 4*1.008 + 1*16.0

= 32.042 g/mol

use:

mass of CH3OH,

m = number of mol * molar mass

= 5.31 mol * 32.04 g/mol

= 170.1 g

Density of water is 1 g/mL or 1 Kg/L

So,

Mass of H2O = density * volume

= 1 Kg/L * 3.21 L

= 3.21 Kg

= 3210 g

Use:

Mass % of CH3OH = Mass of CH3OH * 100 / mass of solution

= 170.1 * 100 / (3210 + 170.1)

= 5.03 %

Answer: 5.03 %